liah0920:

At the beginning of the week, Naveen has 87 trading cards. On Monday, he gets 9 more. On Tuesday, he gives 16 away. On Wednesday, he is given 3 by a friend. On Thursday, he sells all of his cards to a collector. The collector pays him a rate of c for every card that Naveen has. If Naveen earns $124.50 from the sale, what is c, the rate that the collector pays for each card?

jhonyy9:

total 87 cards Monday get 9 so 87 +9 = 96 Tuesday gives 16 away so 96-16=80 Wednesday given 3 so 80+3 = 83 Thursday sells all cards with a rate of ,,c" for every card Naveen earns $124,50 from the sale ,,c" = ? so above step by step we wrote how many cards has Naveen on Thursday ,what is 83 cards and we know that he earns total $124,50 and using these details we can write this equation of 83 * c = 124,50 - from this above wrote equation you need calcule the value of,,c" hope helped understandably easy

Ruthy:

87+9-16+3=83,gave all away for12450 Then 12450/83= 150. Therefore the man paid 150 for each card

jhonyy9:

no bc, you missed the comma - there is 124,50 not 12450